Identify X in the reaction below.
X + 1n → 142Ba + 91Kr + 3 1n
235Th
235U
238U
234Pa
1
Expert's answer
2018-05-28T07:00:38-0400
Solution:
A = 142 + 91 + 3 x 1 – 1 = 235
Z = 56 + 36 + 3 x 0 – 0 = 92
X = 92235U
Answer:
235U